please answer within the next 30 minutes!!! thank you!

consider the following data on an asset. Cost of asset I=$1,000. Useful life, N=5 years, Salvage value, S=$80, what is the depreciation amount in year 3 using the double-declining balance (DBB) depreciation method(multiplier 2)?
Answer #1

Depreciation Rate=(1/useful life)*100

=(1/5)*100

=20%

double declining formula= 2*Cost of assets * Depreciation Rate

so, Double declining Rate will be 2* 20% = 40%.

now year 1 Depreciation = 1000*40%

= $400

now Book Value for year 2 is = 1000-400 = 600.

now year 2 depreciation= 600*40%

=$240.

now book value for year 3 is = 600-240 =360.

now year 3 depreciation= 360*40%

=$144.

so in year 3 Depreciation amount is $144.